OpenStudy (anonymous):

A weight lifter lifts 400 lbs on his fifth attempt, thereby increasing his average by 5 lbs. What is his average after the fifth attempt?

OpenStudy (anonymous):

The average μ for his first 4 attempts is: μ = (x + x + x + x)/4 = 4x/4 = x = μ The average after the fifth lift is (x + x + x + x + 400)/5 = (4x + 400)/5 = x + 5 = μ + 5 So obviously Sx after the 5th lift was 25 lb more than Sx after the first 4. So the average for the first 4 was 375 lb each. And his average after the fifth attempt is 375 x 4 =1500 400 x 1= ..400 ---------------------- So μ = 1900/5 = 380 lb. Answer: The average after the fifth attempt was 380 lb